what are the roots if in ractored form it is y=(2x-3)(3x+4)?
what is the axis of symmetry?

Found 2 solutions by AnlytcPhil, twilightlover:
Answer by AnlytcPhil(1806)

>>what are the roots if in factored form it is y=(2x-3)(3x+4)?<<
----------------------------
The "roots" are found by replacing y by 0 and solving for x
y = (2x-3)(3x+4)
0 = (2x-3)(3x+4)
set each parenthetical factor = 0
2x-3=0, or x = 3/2
3x+4=0, or x = -4/3
-----------------------------
what is the axis of symmetry?
The equation of the axis of symmetry is
x = (the average of the two roots)
or rather:
x = (r1 + r2)/2
x = [(3/2) + (-4/3)]/2
x = (3/2 - 4/3)/2
x = (9/6 - 8/6)/2
x = (1/6)/2
x = 1/12
That's its equation, x = 1/12
Edwin

Answer by twilightlover(1)

(2x-3)
2x=+3
x=3/4
(3x+4)
3x=-4
x=-4/3